OpenStudy (mackenzie2013):

How many times will 1/14 go into 6 7/10?

OpenStudy (mackenzie2013):

THe only options I have is 6 1/14 6 1/20 93 4/5 93 7/10

Directrix (directrix):

>>670/14 is not correct ----------------- (67/10) / (1/14) = (67/10) *(14/1) = 938 / 10 = ? @Mackenzie2013

OpenStudy (mackenzie2013):

So 93 4/5?

Directrix (directrix):

Yes. 93.8 is the same as 93 4/5

OpenStudy (anonymous):

Don't know where I went wrong. Sorry.